OpenStudy (anonymous):

The sum of two numbers is 81. The difference of twice one and three times the other is 62. Find the numbers. How do i do this problem?!?!?!

OpenStudy (anonymous):

Using Systems of linear equations but how?!?!

OpenStudy (anonymous):

uhhhhhhhhhhhhhhhhhhhhhhhh

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

try: let x = the first number let y = the second number Then follow the statements to make your equation. Let me know if you need more hints

OpenStudy (jdoe0001):

so let's call this 2 numbers "a" and "b" their sum is 81 => a + b = 81 twice one and three times the other is 62. 2a 3b = 62 so a + b = 81 2a + 3b = 62 so, there :)

OpenStudy (jdoe0001):

ohh.. the difference... shoot...ok

OpenStudy (jdoe0001):

so let's call this 2 numbers "a" and "b" their sum is 81 => a + b = 81 twice one and three times the other is 62. 2a - 3b = 62 so a + b = 81 2a - 3b = 62 so, there :)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."